# Simulating fermions with exponentially lower overhead

### v1

@@ Abstract
Simulating time evolution under fermionic Hamiltonians is a compelling application of quantum computers because it lies at the core of predicting the properties of materials and molecules. Fermions can be simulated on qubit-based quantum computers using a fermion-to-qubit mapping, subject to an overhead—the circuit depth on a qubit quantum computer divided by that on a quantum computer built from native fermionic modes—at worst scaling linearly with the number of modes $N$. Existing approaches that lower this depth overhead usually trade it for space, using $O(N)$ ancilla qubits. We exponentially reduce the worst-case overhead of ancilla-free fermion-to-qubit mappings to $O(\log^{2}N)$ by constructing circuits that perform any fermionic permutation on qubits in the Jordan-Wigner encoding in depth $O(\log^{2}N)$. We also show that our result generalizes to permutations in any product-preserving ternary tree fermionic encoding. When introducing $O(N)$ ancillas and mid-circuit measurement and feedforward, the overhead reduces to $O(\log N)$. Finally, we show that our scheme can be used to implement the fermionic fast Fourier transform, a key subroutine in chemistry simulation, with overhead $\Theta(\log N)$ without ancillas and $\Theta(1)$ with ancillas, improving exponentially over the best previously known ancilla-free algorithm with overhead scaling linearly with $N$. Our results show that simulating fermions with qubit quantum computers comes at a much lower asymptotic overhead than previously thought.
@@ I Introduction
Simulating fermionic Hamiltonians is one of the most promising applications of quantum computing, enabling the solution of classically challenging problems in quantum chemistry [1], materials science [2], and high-energy physics [3]. However, in order to simulate fermions with qubit quantum computers, the fermionic operators must be mapped to qubit operators. This leads to a muliplicative overhead in the circuit depth and gate count. In other words, we define this overhead as the depth or gate count on a conventional qubit quantum computer (see e.g. Refs. [4, 5, 6, 7] for recent experimental demonstrations) divided by the depth or gate count on a fermionic quantum computer that digitally manipulates fermionic particles directly (see e.g. Refs. [8, 9, 10, 11]).
The simplest mapping of fermions to qubits is the Jordan-Wigner encoding [12]. This encoding has depth overhead of $O(N)$ for $N$ fermionic sites because some fermionic Hamiltonian terms are mapped to Pauli strings with support on $\Theta(N)$ sites, even if in the fermionic formulation these terms only have support on $\Theta(1)$ sites. This is because the Jordan-Wigner encoding labels the fermionic modes along a fictitious one-dimensional chain. For example, the hopping term from one end of the chain to the other has weight $2$ in the fermionic formulation and weight $N$ after mapping to qubits. This is a highly non-local hopping term, but because of the effective one-dimensionality, even geometrically local hopping terms in low dimension can have high weight if they are not along the direction of the Jordan-Wigner chain. For example, the hopping term in two dimensions which is not along the chain has weight $\Theta(\sqrt{N})$ [5]. While it is widely known that this leads to an at-worst linear depth overhead for fermion time evolution, the reason for this is often miscommunicated: it is not the high weight of the fermion terms which leads to the depth overhead, but the restriction in parallelization. The most straightforward approach for implementing time evolution under such a Hamiltonian is done with a so-called Pauli gadget, i.e., the conjugation of a single-qubit rotation with an $N$-qubit CNOT ladder, which requires depth $N$. However, this CNOT ladder can actually be compressed to depth $O(\log N)$ without ancillas [13]. Therefore, a single fermionic Hamiltonian term can be implemented in depth $O(\log N)$ within the Jordan-Wigner encoding, and therefore the depth overhead resulting from the $O(N)$ operator weight is only $O(\log N)$. Instead, the reason why the Jordan-Wigner encoding has $O(N)$ depth overhead for the overall time-evolution task is the parallelization restriction: in the worst case, only one Hamiltonian term can be applied at a time, resulting in a depth scaling with the number of Hamiltonian terms, which is $\Omega(N)$ for a non-trivial Hamiltonian. By contrast, for fermionic quantum computers, there is no such restriction other than the connectivity graph of the Hamiltonian—in most physical Hamiltonians, the connectivity is such that $\Theta(N)$ terms can be implemented in parallel (this is for example the case in the 2D Fermi-Hubbard model [14]), such that for example $\Theta(N)$ Hamiltonian terms can be implemented with depth $\Theta(1)$.
The Bravyi-Kitaev encoding [15] might at first seem to address this issue [16], as it reduces the operator weight after the fermion-to-qubit mapping to only $O(\log N)$ sites. However, unfortunately, these savings for a single fermionic Hamiltonian term do not carry over to the task of implementing all terms: the Pauli strings to which the fermion creation and annihilation operators are mapped all have one qubit in common, so they cannot be implemented in parallel [17]. Therefore, in practice, even though the weight reduction can lead to a gate count improvement, for most problems the Bravyi-Kitaev encoding does not show an improvement over the $O(N)$ depth overhead of the Jordan-Wigner encoding. Schemes have been proposed to avoid this restriction by using ancillas [15, 18, 17, 19, 10], which in particular can achieve $O(1)$ overhead for geometrically local models [20, 21]. However, all of these approaches require $\Omega(N)$ ancillas, which might be a prohibitive overhead, especially for fault-tolerant architectures, where qubit count is often more expensive than depth due to the exponential error suppression [22].
In this work, we show that arbitrary fermionic models (which are fermion parity-conserving, i.e., with terms that are products of an even number of creation/annihilation operators, but without any restrictions on spatial locality or $k$-locality) can be time-evolved with a worst-case depth overhead of only $O(\log^{2}N)$ in the Jordan-Wigner encoding, exponentially improving the best previous bound $O(N)$ on the overhead. Crucially, we do not require any ancillas. When introducing O(N){O}(N) ancillas as well as measurement and feedforward, the depth overhead reduces to $O(\log N)$. In addition, we show how any product-preserving ternary tree encoding (Definitions 5 and 7), such as the Bravyi-Kitaev encoding, can be mapped in depth $O(\log^{2}N)$ to the Jordan-Wigner encoding, implying that fermionic time evolution for any encoding in this class can be done with depth overhead $O(\log^{2}N)$. Finally, we apply our methods to the problem of the fermionic fast Fourier transform (FFFT), an important subroutine in materials simulation [2]. We show that the FFFT can be applied in depth $O(\log^{2}N)$, which is only a factor of $O(\log N)$ more than with fermionic quantum computers [20, 10], and exponentially better than the best previous ancilla-free scheme that required worst-case depth linear in $N$ [23]. Combining this result with those in Ref. [24] (which use $\Theta(N\log N)$ ancillas to implement Coulomb interactions in log-depth) also immediately implies that a single Trotter step of the quantum chemistry Hamiltonian in the plane-wave basis can be implemented in polylog depth by a quantum computer using only $\tilde{O}(N)$ qubits, exponentially improving over the best previous Trotter step for this problem with $\tilde{O}(N)$ qubit count [2]. A previous result existed which achieved polylog-depth [10] with $O(N^{2})$ qubits—our result reduces the qubit count to $\tilde{O}(N)$. For the end-to-end complexity of time-evolving under the second-quantised quantum-chemistry Hamiltonian in the plane wave basis with $N$ plane waves up to time $T$, our results imply that this task can be achieved in depth $\tilde{O}(NT)$ with $\tilde{O}(N)$ qubits, which is the first time this almost-linear scaling has been achieved (we discuss these results in detail in Section V and compare to existing results in Table 1).
Our results build on a reformulation of time evolution in the Jordan-Wigner encoding using fSWAP networks [23] that was shown to be optimal for time-evolving fully connected fermionic models [25]. FSWAP networks solve the parallelization issue for all-to-all connected models by changing the ordering of the fermionic modes in the Jordan-Wigner encoding by nearest-neighbour transpositions of modes. However, it was not previously clear how to generalize this to less connected models. We show that a compilation of fSWAP networks using intuition gained from neutral-atom qubit routing [26] enables permuting fermions arbitrarily in-between Trotter layers using a circuit of only $O(\log^{2}N)$ depth. After permutation, the modes are arranged such that the same parallelization is offered as in fermionic quantum computers for arbitrary fermionic models. Our results therefore suggest that the Jordan-Wigner encoding may be close to optimal in both qubit count and circuit depth, at least asymptotically and without considering any details of the Hamiltonian (such as spatial locality in $2$ or $3$ dimensions, in which case the overhead can be reduced to $\Theta(1)$ [21]).
Problem statement and main result
As we will discuss near the end of this section, both of these restrictions are not necessary: our scheme generalizes to any Trotter scheme and any Hamiltonian (without any restrictions on $k$-locality).
On a fermionic quantum computer, which in particular can directly implement $\exp(-itJ_{ij}(a^{\dagger}_{i}a_{j}+\mathrm{h.c.}))$ directly as a single gate, the minimal depth to implement this scheme depends only on how many of the $J_{ij}$ are nonzero and to what extent the structure of the $J_{ij}$ may be parallelized. For instance, for a full-rank tensor, i.e., with $\Theta(N^{2})$ non-zero terms, the terms can be grouped such that $O(N)$ can be implemented in parallel. This means that the minimal depth is $\Omega(N)$. For a non-full-rank tensor, the situation is more complicated: if there are only $\Theta(N)$ non-zero terms, this may still require depth $\Theta(N)$ if all of the terms have one fermion in common. In order to discuss the overhead introduced by the fermion-to-qubit encoding independently of the parallelization restrictions introduced by the Hamiltonian parameters, we define the fermion-to-qubit overhead as the multiplicative overhead in gate depth between an implementation on a fermionic quantum computer and a qubit quantum computer, or in other words, the gate depth on a qubit quantum computer divided by the gate depth on a fermion quantum computer. To make this comparison, we assume all-to-all connectivity in both qubit and fermion quantum computers.
In this work, we show that fSWAP networks can in fact achieve a worst-case upper bound of $O(\log^{2}N)$ depth overhead compared to fermionic quantum computers for general fermionic Hamiltonians. We do so by generalizing the idea of fermion transpositions in fSWAP networks [23] to arbitrary-range fermionic permutations. Specifically, we split the time evolution into alternations of arbitrary, efficient fermion permutations and time evolutions. For our example, to order $t^{2}$,
The terms are grouped into $N_{\mathrm{layers}}$ layers, so that terms within a given layer can be applied in parallel on a fermionic quantum computer. The goal of the permutations is to rearrange the order of the fermion sites in a given layer so that the Hamiltonian terms only act on sites that are adjacent in the Jordan-Wigner encoding, such that $k$-local fermion terms map to $k$-local qubit terms. More specifically, we define the product of operators using the following order: $\prod_{i=1}^{N}O_{i}=O_{N}\cdots O_{2}O_{1}$. In the first line of Eq. (2), we insert resolutions of identity in terms of permutation unitaries $U_{\sigma_{l}}$, which apply permutation $\sigma_{l}$ corresponding to layer $l$: $U_{\sigma_{l}}a_{i}U_{\sigma_{l}}^{-1}=a_{\sigma_{l}(i)}$. To enable parallelization of all terms within a given layer in the Jordan-Wigner encoding, the permutation unitaries $U_{\sigma_{l}}$ reorder the fermions along the Jordan-Wigner chain such that, for all $J_{ij}$ within a given layer, the fermions $\sigma_{l}(i)$ and $\sigma_{l}(j)$ are neighboring along the Jordan-Wigner chain. This means that $\exp(-itJ_{ij}(a^{\dagger}_{\sigma_{l}(i)}a_{\sigma_{l}(j)}+\mathrm{h.c.}))\rightarrow\exp(-itJ_{ij}(\Sigma^{+}_{\sigma_{l}(i)}\Sigma^{-}_{\sigma_{l}(j)}+\mathrm{h.c.}))$, where $\Sigma^{\pm}=(X\pm iY)/2$ are qubit Pauli operators, can be implemented in depth $O(1)$ and no parallelization restriction arises from the fermion-to-qubit mapping. Furthermore, for every $l<N_{\mathrm{layers}}$, $U_{\sigma_{l+1}}U_{\sigma_{l}}^{-1}$ can be combined into a single permutation unitary.
This construction therefore reduces the fermion-to-qubit overhead to the circuit depth for arbitrary fermion permutations $U_{\sigma_{l}}$. We show an exponentially smaller upper bound than the previously-believed $O(N)$ bound on this task:
@@
We show this theorem by explicitly constructing the circuit, also enabling a practical implementation. In addition, we show the following corollary:
@@
For any product-preserving ternary tree encoding of fermions into $N$ qubits, there exists a circuit of depth $O(\log N)$ with $O(N)$ ancillas and measurement and feedback that implements any given fermionic permutation of $N$ fermionic modes.
As an application, we show that the fermionic fast Fourier transform [20] can be implemented in depth $O(\log^{2}N)$ without ancillas and $O(\log N)$ with ancillas.
The remainder of this manuscript is organized as follows. In Section II, we introduce definitions and notation. In Section III, we present a $O(\log^{2}N)$-depth circuit that performs any fermionic permutation under the Jordan-Wigner mapping. In Section IV, we present a $O(\log^{2}N)$-depth circuit that converts between any two product-preserving ternary tree mappings. In Section VI, we briefly discuss some directions for future work.
@@ III Permutations in Jordan-Wigner
Next, we introduce a technical circuit construction which we use both here and in Sec. IV.
@@
We begin by decomposing $\sigma$ into $O(\log N)$ layers of disjoint staircase permutations using Lemma 1. Since, in each layer, each staircase permutation has a disjoint range of elements, by promoting each of the transpositions in the $\lceil\log_{2}N\rceil$ layers of staircase permutations into fermionic swaps in the Jordan-Wigner encoding, one obtains a fermionic routing circuit similarly made of $\lceil\log_{2}N\rceil$ layers of fermionic staircase permutations acting on disjoint sets of qubits. This is seen by the property that each $\operatorname{fSWAP}_{ij}$ acts only on qubits $\{i,i+1,\dots,j\}$, so each fermionic staircase permutation acts on disjoint sets of qubits in each layer, and thus the problem of fermionic routing is reduced to implementing arbitrary fermionic staircase permutations, which we now show can be done in $O(\log N)$ depth.
@@ IV Transformations between Encodings
In this section, we extend the regime of fermion-to-qubit mappings for which the fermion permutation circuit of Theorem 2 is applicable. We show that a broad range of mappings, called product-preserving ternary tree mappings, can be transformed into the Jordan-Wigner mapping in depth $O(\log^{2}N)$. Consequently, routing in those mappings can also be implemented in depth $O(\log^{2}N)$ by first transforming to Jordan-Wigner, applying our routing circuit, and then transforming back to the original map.
Let $\phi_{1}$ and $\phi_{2}$ be two fermion-to-qubit maps over the same fermion and qubit spaces. Since we assume $\dim\mathcal{H}_{f}=\dim\mathcal{H}_{q}=2^{N}$, they are unitary maps. Then there exists a unitary $U\in\mathcal{B}(\mathcal{H}_{q})$ transforming between the two maps such that $\phi_{2}=U\phi_{1}$, namely $U=\phi_{2}\phi_{1}^{-1}$. This transformation acts on operators via conjugation:
@@
We construct efficient circuits implementing $U$ when $\phi_{1}$ and $\phi_{2}$ lie in a class of fermion-to-qubit maps known as product-preserving ternary tree maps.
Ternary tree maps were introduced in Ref. [29] to give a mapping that provably minimizes the Pauli weight of $\Phi(a_{k})$ averaged over all $k$. However, the initial construction only focused on the mapping of operators and not of states. Subsequent works [32, 33] extended the construction to study properties of the mapping of states. In this paper, we use a formulation of ternary trees adapted from Ref. [34]. See any standard classical algorithms text (e.g. Ref. [35]) for the basics of trees and tree traversal algorithms.
@@
[…]
(a) Jordan-Wigner
[…]
(b) Parity
[…]
(c) Bravyi-Kitaev
Figure 11: Many common fermion-to-qubit mappings, including (a) Jordan-Wigner, (b) Parity, and (c) Bravyi-Kitaev, can be expressed as ternary trees. All three examples shown are binary-shaped ternary tree maps.
@@
The Jordan-Wigner, Bravyi-Kitaev, and Parity (defined as $\phi_{P}\left|x_{0},\dots,x_{N-1}\right\rangle_{f}=\left|p_{0},\dots,p_{N-1}\right\rangle$ with $p_{k}$ given by Eq. (10)) maps can all be represented as binary-shaped ternary tree maps [32], as shown in Fig. 11. Furthermore, in all three examples, the qubits are always labeled via an inorder traversal, and the leaves are labeled in order from left to right. (Recall that an inorder traversal is an ordering of nodes where we first recursively perform an inorder traversal of the left subtree, then visit the root, and then recursively perform an inorder traversal of the right subtree.)
@@
Figure 12: Tree rotation corresponding to the adjoint action of the CNOTjk transformation. The square boxes can be leaves or contain further subtrees.
@@
We remark that, while Theorem 3 was formulated in the context of fermion-to-qubit maps, the circuits can also transform qubit states in the same manner, outside the context of any fermions. In particular, the same circuit constructed in Theorem 3 to transform Jordan-Wigner to Parity also gives a log-depth circuit implementing the parity transform [Eq. (10)] on qubit states. We also remark that this reproduces the same circuit as the well-known divide-and-conquer $\operatorname{CNOT}$ circuit for parity.
For the next transformation, we use the following result regarding compressing $\operatorname{CNOT}$ circuits.
@@
Let $\mathcal{C}=U_{1}U_{2}\dots U_{n}$, where each $U_{k}$ is either $\operatorname{CNOT}_{q_{0},q_{k}}$ or $\operatorname{CNOT}_{q_{k},q_{0}}$ and $q_{0},q_{1},\dots,q_{n}\in[N]$ are all distinct. Then $\mathcal{C}$ can be implemented by a circuit of depth $O(\log n)$.
@@
Without loss of generality, we can relabel the qubit indices so that $q_{k}=k$ for each $0\leq k\leq n$. First, we apply a Hadamard transform to qubit 0 using the following identities:
$H_{0}\operatorname{CNOT}_{0,k}H_{0}$ $=H_{k}\operatorname{CNOT}_{k,0}H_{k},$ (15a)
$H_{0}\operatorname{CNOT}_{k,0}H_{0}$ $=\operatorname{CZ}_{k,0}.$ (15b)
In particular, $H_{0}U_{k}H_{0}$ still only acts on qubits 0 and $k$, so any new $H_{k}$ can be commuted to the ends of the circuit, as depicted in Fig. 14. Next, we commute all the $\operatorname{CZ}$ gates to the left. Let $S$ be the set of indices $j$ such that $U_{j}=\operatorname{CNOT}_{j,0}$. For $j<k$, we have
$\operatorname{CZ}_{k,0}\operatorname{CNOT}_{j,0}=\operatorname{CNOT}_{j,0}\operatorname{CZ}_{j,k}\operatorname{CZ}_{k,0},$ (16)
so commuting $\operatorname{CZ}_{k,0}$ to the left picks up a $\operatorname{CZ}_{j,k}$ gate for every $j\in S$, giving rise to a $\operatorname{CZ}$-fanout with control $k$ and targets in $S$. Commuting all $\operatorname{CZ}_{k,0}$ to the left, we obtain a contiguous-range $\operatorname{CZ}$-fanout circuit, as shown in Fig. 14, which we showed in Lemma 2 can be implemented in depth $O(\log n)$. Finally, the remaining $\operatorname{CNOT}_{k,0}$ form a $\operatorname{MOD}_{2}$ gate, which can also be implemented in depth $O(\log n)$. ∎
Figure 14: The $\operatorname{CNOT}$ ladder of Lemma 4 is equivalent to Hadamards conjugating a contiguous-range $\operatorname{CZ}$-fanout and a parity gate, each of which can be implemented in log depth.
@@
There exists a $\operatorname{CNOT}$ circuit of depth $O(\log^{2}N)$ transforming between any two binary-shaped ternary trees on $N$ qubits sharing the same ordering of Majorana leaf labeling.
@@
We show that there exists a circuit of depth $O(\log^{2}N)$ transforming any such binary-shaped tree to a characteristic binary tree on $N$ nodes. We choose this characteristic binary tree to be one with the property that for every node, the number of nodes in the left subtree minus the number of nodes in the right subtree is either 0 or $1$ (i.e., the tree is weight-balanced favoring a heavy left). Such a tree exists for every $N$ and can be constructed inductively.
The transformation algorithm is provided in Algorithm 1. First, perform an inorder traversal on nodes to locate the node $r$ that should become the root. Because the tree is weight-balanced favoring a heavy left, this will always be the $\left\lceil\frac{N+1}{2}\right\rceilth$ node traversed. Then, let $q_{1},\dots,q_{h}=r$ be the path of nodes from the root to $r$. We rotate $r$ up to the root by applying $\operatorname{CNOT}$ gates between qubits $(r,q_{h-1}),\dots,(r,q_{1})$, each oriented so that node $r$ is rotated upward. This step is depicted in Fig. 15. By Lemma 4, this can be implemented in depth $O(\log h)$. As $h\leq N$, this is $O(\log N)$.
Finally, we recurse on the left and right subtrees. The subtrees are disjoint, so they can be transformed in parallel, and each subtree will have at most half the size. Thus the recursion terminates in $O(\log N)$ iterations, each of which has depth $O(\log N)$, for a total depth of $O(\log^{2}N)$.
At the end of this algorithm, we will have a binary-shaped tree with the characteristic shape. By the inorder invariant of tree rotations, the Majorana leaves remain labeled in the same order. The only remaining degree of freedom is in the qubit labels, which can be fixed in depth two by decomposing the desired qubit permutation into two layers of $\operatorname{SWAP}$ gates [37]. ∎
Algorithm 1 Circuit transforming any binary subtree into any other binary tree shape

### v2

# Low-depth fermion routing without ancillas

@@ Abstract
Routing is the task of permuting qubits in such a way that quantum operations can be parallelized maximally, given constraints on the hardware geometry. When simulating fermions in the Jordan-Wigner encoding with qubits, a one-dimensional nearest-neighbor-connected geometry is effectively imposed on the system, independently of the underlying hardware, which means that naively, an $O(N)$ depth routing overhead is incurred. Recently, Maskara et al. [arXiv:2509.08898] demonstrated that this routing overhead can be reduced to $O(\log N)$ by decomposing general fermion routing into $O(\log N)$ interleave permutations of depth $O(1)$, using $\Theta(N)$ ancillary qubits and employing measurements and feedforward. Here, we exhibit an alternative construction that achieves the same asymptotic performance. We also generalize the result in two ways. Firstly, we show that fermion routing can be performed in depth $O(\log^{2}N)$ without ancillas, measurements, or feedforward. Secondly, we construct efficient mappings with $O(\log^{2}N)$ depth between all product-preserving ternary tree fermionic encodings, thereby showing that fermion routing in any such encoding can be done efficiently. While these results assume all-to-all connectivity, they also imply upper bounds for fermion routing in devices with limited connectivity by multiplying the fermion routing depth by the worst-case qubit routing depth.
@@ I Introduction
Standard quantum computers based on qubit architectures suffer from locality constraints. These constraints determine which pairs of qubits are allowed to interact, and which interactions may be implemented in parallel. Hence, these constraints present a challenge for performing quantum computation, and must be accounted for in order to implement quantum information processing tasks. One way to handle locality constraints is by performing routing, which is the task of implementing permutations of qubits. By permuting qubits, one may simulate all-to-all connectivity and therefore implement arbitrary quantum circuits with an overhead in circuit depth. In order to minimize this overhead, the problem of routing efficiently under interaction constraints has been studied using swap gates [1, 2, 3], teleportation and mid-circuit measurements [4, 5, 6], and dynamic reconfigurability [7, 8]. In particular, the worst-case depth of routing of qubits arranged in a one-dimensional line using reconfigurable neutral-atom hardware has been shown to be $O(\log N)$  [9, 8].
[…]
More specifically, we show the following theorem:
@@
We show this theorem by explicitly constructing the circuit, also enabling a practical implementation. In addition, we show the following corollary, using Ref. [11]:
@@
For any product-preserving ternary tree encoding of fermions into $N$ qubits, there exists a circuit of depth $O(\log N)$ with $O(N)$ ancillas and measurement and feedforward that implements any given fermionic permutation of $N$ fermionic modes.
The remainder of this manuscript is organized as follows. In Section II, we introduce definitions and notation. In Section III, we present a $O(\log^{2}N)$-depth circuit using staircase permutations that performs any fermionic permutation under the Jordan-Wigner mapping. In Section IV, we present a $O(\log^{2}N)$-depth circuit that converts between any two product-preserving ternary tree mappings. In Section V, we discuss applications to quantum simulation of fermionic models. In Section VI, we briefly discuss some directions for future work.
@@ III Permutations in Jordan-Wigner
We remark that these two stages resemble the construction in Ref. [11], where an arbitrary permutation is decomposed into $O(\log N)$ layers of special efficient permutations. However, we differ in the choice of these special efficient permutations: we use ones that we call staircase permutations, while Ref. [11] uses others that they call interleaves. While some permutations are both interleaves and staircase permutations, neither is a subset of the other. Every staircase permutation can be implemented as a product of two interleaves [8, Sec. SI], and it remains open whether every interleave can be implemented via a constant number of staircase permutations.
@@
Next, we introduce a technical circuit construction that is equivalent to [11, Lemma 3] but that we rederive here to demonstrate ancilla-free techniques.
@@
We begin by decomposing $\sigma$ into $O(\log N)$ layers of disjoint staircase permutations using Lemma 1. Since, in each layer, each staircase permutation has a disjoint range of elements, by promoting each of the transpositions in the $\lceil\log_{2}N\rceil$ layers of staircase permutations into fermionic swaps in the Jordan-Wigner encoding, one obtains a fermion routing circuit similarly made of $\lceil\log_{2}N\rceil$ layers of fermionic staircase permutations acting on disjoint sets of qubits. This is seen by the property that each $\operatorname{fSWAP}_{ij}$ acts only on qubits $\{i,i+1,\dots,j\}$, so each fermionic staircase permutation acts on disjoint sets of qubits in each layer, and thus the problem of fermion routing is reduced to implementing arbitrary fermionic staircase permutations, which we now show can be done in $O(\log N)$ depth.
@@ IV Transformations between Encodings
In this section, we extend the regime of fermion-to-qubit mappings for which the fermion permutation circuit of Ref. [11] and Theorem 2 is applicable. We show that a broad range of mappings, called product-preserving ternary tree mappings, can be transformed into the Jordan-Wigner mapping in depth $O(\log^{2}N)$. Consequently, routing in those mappings can also be implemented in depth $O(\log^{2}N)$ by first transforming to Jordan-Wigner, applying our routing circuit, and then transforming back to the original map.
Let $\phi_{1}$ and $\phi_{2}$ be two fermion-to-qubit mappings over the same fermion and qubit spaces. Since we assume $\dim\mathcal{H}_{f}=\dim\mathcal{H}_{q}=2^{N}$, they are unitary maps. Then there exists a unitary $U\in\mathcal{B}(\mathcal{H}_{q})$ transforming between the two mappings such that $\phi_{2}=U\phi_{1}$, namely $U=\phi_{2}\phi_{1}^{-1}$. This transformation acts on operators via conjugation:
@@
We construct efficient circuits implementing $U$ when $\phi_{1}$ and $\phi_{2}$ lie in a class of fermion-to-qubit mappings known as product-preserving ternary tree mappings.
Ternary tree mappings, also known as qubit trees in other contexts [23], were introduced in Ref. [19] to give a mapping that provably minimizes the Pauli weight of $\Phi(a_{k})$ averaged over all $k$. However, the initial construction only focused on the mapping of operators and not of states. Subsequent works [24, 25] extended the construction to study properties of the mapping of states. In this paper, we use a formulation of ternary trees adapted from Ref. [26]. See any standard classical algorithms text (e.g. Ref. [27]) for the basics of trees and tree traversal algorithms.
@@
Figure 11: Many common fermion-to-qubit mappings, including (a) Jordan-Wigner, (b) Parity, and (c) Bravyi-Kitaev, can be expressed as ternary trees. All three examples shown are binary-shaped ternary tree mappings.
@@
The Jordan-Wigner, Bravyi-Kitaev, and Parity (defined as $\phi_{P}\left|x_{0},\dots,x_{N-1}\right\rangle_{f}=\left|p_{0},\dots,p_{N-1}\right\rangle$ with $p_{k}$ given by Eq. (8)) mappings can all be represented as binary-shaped ternary tree mappings [24], as shown in Fig. 11. Furthermore, in all three examples, the qubits are always labeled via an inorder traversal, and the leaves are labeled in order from left to right. (Recall that an inorder traversal is an ordering of nodes where we first recursively perform an inorder traversal of the left subtree, then visit the root, and then recursively perform an inorder traversal of the right subtree.)
@@
Figure 12: Tree rotation corresponding to the adjoint action of the CNOTjk transformation. The triangles can be leaves or contain further subtrees.
@@
We remark that, while Theorem 3 was formulated in the context of fermion-to-qubit mappings, the circuits can also transform qubit states in the same manner, outside the context of any fermions. In particular, the same circuit constructed in Theorem 3 to transform Jordan-Wigner to Parity also gives a log-depth circuit implementing the parity transform [Eq. (8)] on qubit states. We note that this reproduces the same circuit as the well-known divide-and-conquer $\operatorname{CNOT}$ circuit for parity.
Algorithm 1 Circuit transforming any binary-shaped ternary tree into the Jordan-Wigner shape
1: while tree is not Jordan-Wigner shape do
2:   for all nodes $k$ on the right spine do
3:    if $k$ has a non-leaf left child $j$ then
4:      Apply $\operatorname{CNOT}_{jk}$
We also remark that the technique of rotating nodes onto the right spine is not restricted to starting from Bravyi-Kitaev. Rather, this technique can be applied to transform any binary-shaped ternary tree to the Jordan-Wigner shape and is detailed in Algorithm 1.
@@
Let $T$ be a binary-shaped ternary tree mapping and let $d$ be the furthest distance from any node to the right spine. There exists a $\operatorname{CNOT}$ circuit of depth $d$ that transforms between $T$ and the Jordan-Wigner tree shape.
@@
The circuit is constructed by Algorithm 1. For each node not already on the right spine, its distance to the right spine decreases by $1$ in each iteration of the while loop. Therefore the while loop runs $d$ times. Within each iteration, all $\operatorname{CNOT}_{jk}$ gates can be applied in parallel as they act on disjoint qubits. Therefore the resulting circuit has depth $d$. ∎
@@
There exists a $\operatorname{CNOT}$ circuit of depth $O(\log N)$ transforming between any two binary-shaped ternary trees on $N$ qubits sharing the same ordering of Majorana leaf labeling.
@@
It suffices to show that there exists a circuit of depth $O(\log N)$ transforming any binary-shaped tree to the Jordan-Wigner tree shape. To do so, we first use parallel tree rotations to transform an arbitrary binary tree into a binary tree with depth $O(\log N)$. The depth upper bounds the maximum distance of any node from the right spine, so by Lemma 4 we can transform the log-depth tree into the Jordan-Wigner shape in depth $O(\log N)$. By the inorder invariant of tree rotations, the Majorana leaves remain labeled in the same order. The only remaining degree of freedom is in the qubit labels, which can be fixed in depth two by decomposing the desired qubit permutation into two layers of $\operatorname{SWAP}$ gates [29].
To attain a log-depth binary-shaped tree, we introduce some notation to identify the imbalances in the tree. The weight of a node $A$, denoted $\operatorname{wt}(A)$, is the number of leaves in the subtree rooted at $A$. The halving depth of $A$, denoted $\operatorname{hd}(A)$, is the minimum number of levels below $A$ such that every descendant $B$ of $A$ in that level has at most half the weight, i.e., $\operatorname{wt}(B)\leq\tfrac{1}{2}\operatorname{wt}(A)$.
There is clearly at most one node per level with weight $\tfrac{1}{2}\operatorname{wt}(A)$—there is not enough weight in the whole subtree for more than one. In particular, some $B$ exists $\operatorname{hd}(A)-1$ steps below $A$ such that $\operatorname{wt}(B)>\tfrac{1}{2}\operatorname{wt}(A)$. By inclusion, all the ancestors of $B$ up to $A$ are also above the threshold. Inspired by this, we define the heavy path from $A$ to be the sequence of descendants (of length $\operatorname{hd}(A)$) that have weight $\tfrac{1}{2}\operatorname{wt}(A)$.
[…]
$A^{\prime}$ to $A^{\prime\prime}$
left right
AA to $A^{\prime}$ left $\operatorname{CNOT}_{A^{\prime\prime}A^{\prime}}$ (Fig. 12) $\operatorname{CNOT}_{A^{\prime\prime}A}\operatorname{CNOT}_{A^{\prime}A^{\prime\prime}}$ (Fig. 14(a))
right $\operatorname{CNOT}_{AA^{\prime\prime}}\operatorname{CNOT}_{A^{\prime\prime}A^{\prime}}$ (Fig. 14(b)) $\operatorname{CNOT}_{A^{\prime}A^{\prime\prime}}$ (Fig. 12)
Table 1: The choice of tree rotations in Algorithm 2.
The key property is that in all cases, we reduce the length of three steps in a heavy path to two or less.
Figure 14: In Theorem 4, depending on the shape of the heavy path, these rotations may be applied. In both cases, $A^{\prime\prime\prime}$ can be the root of either subtree $c$, $d$ or $e$.
Now consider how a round of the algorithm affects the halving depth $d=\operatorname{hd}(X)$ of an arbitrary node $X$. The heavy path has $d$ nodes, $X=X^{(1)}$ through $X^{(d)}$. The last 33 ($X^{(d-2)}$, $X^{(d-1)}$, $X^{(d)}$) are too close to the end to trigger a tree rotation in the algorithm. Of the remaining nodes, at least $\lfloor\tfrac{d}{3}\rfloor-1$ appear on a level that is a multiple of three.
The tree rotations usually change the root of the subtree; call the new root $Y$. The final element of the original heavy path, $X^{(d)}$, exists in the subtree and still has weight $\operatorname{wt}(X^{(d)})>\tfrac{1}{2}\operatorname{wt}(Y)$. However, the algorithm has applied no tree rotation that could affect $X^{(d)}$, so its children have weight $\leq\tfrac{1}{2}\operatorname{wt}(Y)$. Hence, the halving depth of $Y$ is its distance from $X^{(d)}$. The distance between the subtree’s root and $X^{(d)}$ decreases by $\lfloor\tfrac{d}{3}\rfloor-1\geq\tfrac{d-5}{3}$ from the tree rotations, so
$\operatorname{hd}(Y)\leq d-\frac{d-5}{3}\leq\frac{2d+5}{3}.$
[…]
Algorithm 2 Circuit transforming any binary-shaped ternary tree into the Jordan-Wigner shape